# Quantum Origin of Classical Background Fields from Coherent States

### v1

@@ Abstract
Classical background electromagnetic fields are routinely employed in quantum electrodynamics to describe a wide range of physical situations, from laser-matter interactions to strong-field phenomena. In this work, we present a first-principles formulation that clarifies the quantum origin of such classical background fields in QED by systematically deriving them from coherent states of the electromagnetic field.
Starting from the operator formulation of QED, we show how scattering amplitudes between coherent states naturally lead to an effective description in terms of background fields, while maintaining a clear separation between the coherent laser mode and other quantized photon degrees of freedom. This framework allows one to consistently incorporate effects beyond the fixed background approximation, such as depletion and backreaction, without assuming any particular field strength or intensity regime.
We further demonstrate how the conventional generating functional with a prescribed background field emerges as a limiting case, corresponding to fixed coherent state boundary conditions. The path integral representation is then obtained as a reformulation of the same underlying Heisenberg picture amplitudes, providing a unified view of operator-based and functional approaches.
Our results establish a general and intensity-independent foundation for QED with coherent background fields, within which the standard formulations of strong-field QED arise as well-defined special cases.
@@ 1 Introduction
The use of classical background electromagnetic (EM) fields is a standard and highly successful strategy in quantum electrodynamics (QED). It underlies a wide range of applications, including laser-matter interactions [Salamin2006, Joachain_Kylstra_Potvliege_2011, Raicher2014], external-field problems [Greiner_Muller_Rafelski_1985], and strong-field QED (also known as nonlinear QED) [Ritus1985, Piazza2012, Fedotov2023]. In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations are quantized on top of it. Despite its practical success, this raises a conceptual question: how does a classical background field emerge from the underlying quantum theory of the EM field, and under what conditions is the fixed background approximation justified?
In conventional treatments, this issue is often bypassed by postulating the background field from the outset, leading to the framework of QED in external fields, or the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. Within this framework, scattering amplitudes are computed in the presence of a given background, and techniques such as Volkov solutions [Joachain_Kylstra_Potvliege_2011, Ritus1985, Piazza2012, Fedotov2023, Volkov1936, Itzykson-Zuber1980] and background-dependent generating functionals [Greiner_Muller_Rafelski_1985, Schwinger1951] are employed. However, the background field itself is not dynamical, and its relation to the quantized EM field remains implicit.
A natural candidate for bridging this gap is provided by coherent states of the EM field [Glauber1963a, Glauber1963b, Sudarshan1963]. Coherent states exhibit classical-like properties while remaining fully quantum mechanical, and they have long been recognized as an appropriate description of ideal laser fields [Sargent1974, Loudon_2000]. From this perspective, a classical background field may be viewed not as an external input but as an effective description arising from scattering of coherent states of the quantized EM field, see Ref. [Frantz1965] as an earlier work in that research direction.
In this paper, we develop a first-principles formulation of QED in which classical background fields are systematically derived from coherent states, without assuming any specific field strength or intensity regime. In what follows, we refer to this formulation as QED with coherent background fields (hereafter referred to as coherent-field QED), to emphasize that the classical backgrounds arise from coherent states of the quantized EM field rather than being imposed externally. Our approach starts from the operator formulation of QED in the Schrödinger and Heisenberg pictures [Itzykson-Zuber1980, Sakurai1967], and considers scattering amplitudes between asymptotic coherent states of the EM field, together with arbitrary in- and out-states of other particles. This construction allows us to maintain a clear separation between the coherent laser mode and other quantized photon degrees of freedom, and to consistently discuss effects beyond the fixed background approximation, such as depletion and backreaction [Ilderton-Seipt2018]. We show that the conventional generating functional of QED with a prescribed background field [Greiner_Muller_Rafelski_1985, Schwinger1951] emerges as a limiting case, corresponding to fixed coherent state boundary conditions. Furthermore, we reformulate the same scattering amplitudes in terms of a path integral representation, demonstrating that the operator-based and functional approaches describe the same physics from complementary perspectives.
While our formulation is not restricted to the high-intensity regime [Ritus1985, Piazza2012, Fedotov2023], it naturally reproduces standard strong-field QED as well-defined limiting cases. In this sense, the present work provides a unified and intensity-independent foundation for QED in background fields, clarifying the conceptual status of classical backgrounds and their domain of validity. The present formulation provides a first-principles reinterpretation of the fixed background approximation commonly employed in strong-field QED (i.e., within the coherent-field QED framework in the fixed-background limit), by relating it explicitly to coherent quantum states of the EM field.
@@
This paper is organized as follows. In Sec. 2, we introduce coherent states of EM fields. Section 3 presents the formalism connecting the Schrödinger and Heisenberg pictures. Section 4 discusses the Furry picture and Volkov solutions, while Sec. 5 applies the formalism to the Heisenberg–Euler vacuum and laser depletion effects. Section 6 relates the approach to Becchi–Rouet–Stora–Tyutin (BRST) quantization, and Sec. 7 presents the path integral formulation. We conclude with a summary and outlook in Sec. 8.
@@ 2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition [Gupta1950, Bleuler1950]. Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
$\partial_{\mu}A_{(+)\mathrm{H}}^{\mu}(x)|\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (1)
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}A_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adapt the Gupta–Bleuler condition to a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state [Glauber1963a, Glauber1963b, Sudarshan1963, Sargent1974, Loudon_2000] is specified by the following eigen equation
@@
where, $\hat{\mathbb{I}}$ is the identity operator [Glauber1963b, Loudon_2000, Frantz1965]. In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
@@ 3 Basic formalism
In this section, we clarify how a classical background field is consistently incorporated within the present operator-based formulation, with particular emphasis on the role of the Schrödinger and Heisenberg pictures. The purpose of this section is twofold. First, we demonstrate how a coherent background field can be introduced in the Schrödinger picture through a displacement operator acting on the quantum gauge field, leading to a clean separation between classical and quantum degrees of freedom at the operator level. Second, we show how the time-dependent background field employed in coherent-field QED emerges naturally upon transforming to the Heisenberg picture. This analysis makes explicit that the time dependence of background fields is a picture-dependent notion rather than a fundamental property of the underlying Hamiltonian. This picture change makes explicit whether the time dependence of the background field is carried by the state or by the operators.
@@
In relevant experiments at high-intensity laser facilities [ELI-NP_2018, CoReLS_Compton_2024, Sarri2025], scattering processes in strong-field QED are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background [Greiner_Muller_Rafelski_1985, Ritus1985, Piazza2012, Moortgat-Pick2009, Hartin2018]. It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian.
From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. As a result, the fixed background approximation, though highly effective in practice, appears conceptually ad hoc when viewed from first principles of quantum field theory.
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states [Itzykson-Zuber1980, Sakurai1967]. Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
Although the Hamiltonian governing the full quantum dynamics is time independent as an operator, the explicit time dependence of the background field commonly used in coherent-field QED arises upon transforming to the Heisenberg picture. In this picture, the time evolution of operators induces a spacetime dependence in the expectation value of the gauge field taken in a coherent state. Thus, the classical background field appearing in the Furry picture should be understood as an emergent, picture-dependent object rather than as a fundamental ingredient of the theory.
This observation provides a resolution of the apparent tension between the use of time-dependent background fields and the time-independent nature of the underlying quantum Hamiltonian. It also clarifies the precise sense in which coherent states serve as the bridge between fully quantized gauge fields and their classical counterparts in coherent-field QED. In the following, we make these statements explicit by constructing the displaced Hamiltonian in the Schrödinger picture and then translating the resulting expressions into the Heisenberg picture, where the connection to the standard background-field formulation becomes manifest.
@@ 3.1 Departure from the Schrödinger picture
We consider QED in the presence of a background laser EM field, aiming to reformulate it into coherent-field QED. To this end, we assume that a coherent background laser field is always present. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
@@
corresponding to the background field approximation.
@@
Both $\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ and $\hat{H}_{\mathrm{S}}=\int d^{3}\boldsymbol{x}\hat{\mathcal{H}}_{\mathrm{S}}(\boldsymbol{x})$ are taken to be time-independent in the Schrödinger picture [Itzykson-Zuber1980, Sakurai1967]. Under this preparation, the Schrödinger equation governing the “coherent laser field plus other particles” state is
$i\hbar\partial_{t}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{H}_{\mathrm{S}}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle.$ (12)
Formally,
@@
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) [Gupta1950, Bleuler1950] evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture, the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
@@
with that formal solution
@@
Thus the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$, with the coherent laser field factored out, incorporates the influence of the laser ($\mathcal{A}$) through its explicit appearance in the evolution operator.
This confirms that the Schrödinger picture evolution operator consistently incorporates the coherent background field through its displacement, while maintaining full quantum dynamics for all other degrees of freedom. In this way, the framework provides a rigorous starting point for connecting the Schrödinger picture with the Heisenberg picture in the presence of a coherent laser background.

### v2

# Coherent-state boundary conditions as the first-principles origin of background fields in QED

@@ Abstract
QED formulated in prescribed classical background electromagnetic fields is a standard framework for strong-field and laser–matter interactions. It is usually treated as a theory modified by externally imposed fields, obscuring its precise relation to full QED and, in particular, the role of asymptotic boundary conditions for the gauge field.
Starting from an operator-based formulation, we show that QED with background fields is not a distinct theory but arises as a well-defined boundary-condition limit of full QED, in which the classical background field emerges from coherent-state boundary conditions on the quantized electromagnetic field. In this limit, the conventional generating functional used in calculations with prescribed background fields is recovered naturally, while relaxing the boundary conditions allows depletion and backreaction effects to be incorporated within the same framework. The central new result is a rigorous operator-level proof of the equivalence between the fixed background-field approximation and coherent-state asymptotic boundary conditions—a formulation that, to our knowledge, has not been made explicit in previous approaches.
We further demonstrate that the apparent time dependence of background fields does not originate from an explicitly time-dependent Hamiltonian, but instead reflects the choice of picture—Schrödinger versus Heisenberg—in the underlying quantum theory. Rather than introducing new properties of coherent states, our analysis provides a first-principles reinterpretation of the fixed background-field approximation as a controlled and picture-dependent limit of full QED.
@@ 1 Introduction
Classical background electromagnetic (EM) fields are a central ingredient in quantum electrodynamics (QED), underpinning a wide range of applications from laser–matter interactions to strong-field and external-field problems [32, 21, 29, 14, 31, 4, 6]. In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations of the EM and matter fields are quantized on top of it. In this sense, one works within a formulation of QED in which classical and quantized EM fields coexist. This framework has proven extraordinarily successful in practice, enabling non-perturbative treatments of intense fields and the description of phenomena such as nonlinear Compton scattering [31, 3, 27, 28], Breit–Wheeler pair production [31, 27, 28, 30], and vacuum polarization effects [14, 7].
Despite this success, the conceptual status of the fixed background field approximation remains subtle. In this paper, we use the term “background field approximation” to denote formulations of QED in which the EM field is decomposed into a prescribed classical background and quantized fluctuations, with the background treated as fixed and unaffected by the quantum dynamics. In conventional formulations—often summarized under the heading of QED in external EM fields or the Furry picture [9, 25]—the background field is introduced as an explicit c-number input, frequently endowed with spacetime dependence from the outset. While operationally well defined, this procedure leaves open a fundamental question: how does a classical background EM field arise from the underlying quantum theory of the field, and in what precise sense is the fixed background an approximation?
A natural bridge between fully quantized gauge fields and classical backgrounds is provided by coherent states of the radiation field [13, 12, 37]. Coherent states possess classical-like expectation values while remaining genuine quantum states, and have long been recognized as an appropriate description of idealized laser fields [34, 23]. It is also well known that scattering amplitudes involving coherent states can be reexpressed as amplitudes computed in the presence of background fields, a correspondence discussed in various contexts since early work such as Ref. [8] and in later analyses of backreaction and depletion effects [19]. These observations suggest that the role of background fields in scattering processes may ultimately be tied to the choice of asymptotic boundary conditions for the quantized EM field.
The central claim of this paper is that QED with prescribed classical background fields is not a distinct or modified theory. Rather, the fixed background field approximation can be identified unambiguously as a controlled boundary-condition limit of full QED. We show that QED with prescribed background fields arises naturally as a limiting case of scattering between coherent states of the quantized EM field, corresponding to fixed coherent-state boundary conditions. In this sense, the background field is not a fundamental input, but an emergent classical object whose status and limitations can be precisely identified.
Our formulation is developed at the operator level, starting from the Schrödinger and Heisenberg pictures of QED [20, 33]. This perspective makes it possible to disentangle two conceptually distinct issues that are often conflated in standard treatments. First, the classical background field is identified with the expectation value of the quantized gauge field in a coherent state, subject to appropriate physical-state conditions, here implemented via the Gupta–Bleuler formalism [15, 2] or, equivalently, BRST cohomology [1, 38, 22]. Second, the spacetime dependence of the background field—commonly taken as an external input—is shown to arise from the choice of picture: while the Hamiltonian remains time independent in the Schrödinger picture, the Heisenberg-picture evolution of operators induces a time-dependent background expectation value. In particular, the time dependence of background fields is not introduced ad hoc at the level of the Hamiltonian, but emerges naturally from the underlying quantum dynamics.
Within this framework, the conventional generating functional of QED with a prescribed background field [14, 36] is recovered as a special case corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic and fully quantum-mechanical description of effects beyond the fixed background field approximation, such as depletion and backreaction [19], without double counting of photon degrees of freedom. For clarity, we refer to this operator-based formulation—in which classical background fields arise as coherent-state expectation values subject to fixed coherent-state asymptotic boundary conditions—as coherent-field QED. More recently, coherent states have been employed as a technical ingredient to relate amplitudes on background fields to gravitational amplitudes via the classical double copy [18]. Our aim here is different: rather than exploiting this equivalence as a computational tool, we reinterpret the fixed background field approximation itself as a limiting procedure based on coherent-state boundary conditions. While operator-based approaches to QED with prescribed background fields—most notably the unstable-vacuum formalism of Fradkin, Gitman and Shvartsman [7], and related analyses by Gavrilov and Gitman [11]—demonstrate that coherent states can mimic prescribed classical fields in specially chosen initial states, these works formulate the background as an externally imposed c-number field and treat the coherent-state displacement operator within the interaction picture. Consequently, the origin of the background field is picture-dependent and not tied to a boundary-condition prescription in full QED. In contrast, the present work identifies the formulation of QED with prescribed background fields itself as a well-defined coherent-state asymptotic boundary-condition limit, formulated at the operator level and independent of picture choice. This allows a first-principles reinterpretation of QED with prescribed background fields and clarifies its precise relation to the underlying quantum theory.
@@
This paper is organized as follows. In Sec. 2 we introduce coherent states of the EM field. Section 3 develops the operator-based formalism connecting the Schrödinger and Heisenberg pictures. In Sec. 4 we discuss the relation to the Furry picture and Volkov solutions, while Sec. 5 presents applications to the Heisenberg–Euler effective action and laser depletion effects. Section 6 comments on the relation to BRST quantization, and Sec. 7 gives the corresponding path-integral formulation based on coherent-state boundary conditions. We conclude in Sec. 8.
@@ 2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition [15, 2]. Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
$\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}(x)|\mathrm{Physical}_{\mathrm{H}}\rangle$ $=0,$ (1)
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adapt the Gupta–Bleuler condition to a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state [13, 12, 37, 34, 23] is specified by the following eigen equation
@@
where, $\hat{\mathbb{I}}$ is the identity operator [12, 23, 8]. In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
@@ 3 Basic formalism
The purpose of this section is purely conceptual. In standard formulations of QED with prescribed background fields, the time dependence of the background is introduced directly at the level of a c-number function $\mathcal{A}^{\mu}(x)$, obscuring its origin in the underlying quantum theory. Here we show that, once background fields are reinterpreted as expectation values in coherent states, the apparent time dependence of the background field does not originate from a time-dependent Hamiltonian, but exclusively from the choice of picture.
@@
We formulate the incorporation of a classical background field within an operator-based framework, emphasizing the role of the Schrödinger and Heisenberg pictures. We first construct a coherent background field in the Schrödinger picture via a displacement operator acting on the quantized gauge field, which leads to a clean separation between classical and quantum degrees of freedom at the operator level. We then show how the time-dependent background fields familiar from coherent-field QED arise naturally upon transforming to the Heisenberg picture. In this way, the time dependence of the background field is identified as a picture-dependent feature, rather than as an intrinsic property of the Hamiltonian.
Related operator-based approaches to QED with prescribed background fields using coherent-state boundary conditions were discussed previously, notably in Ref. [7, 11]. In those formulations, asymptotic coherent states are introduced via a displacement operator described as time “independent,” implicitly within an interaction- or asymptotic-picture setting. Here, we make the role of picture dependence explicit by showing how the time dependence of the background field arises from operator evolution in the Heisenberg picture while remaining absent in the Schrödinger picture by construction.
@@
In strong-field QED/nonlinear QED experiments at high-intensity laser facilities [10, 24, 35], scattering processes are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background [14, 31, 4, 25, 16]. It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian. From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field [31, 4, 6, 9, 25]. In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. The detailed construction of the Furry picture in coherent-field QED will be developed in Sect. 4.
@@ 3.1 Departure from the Schrödinger picture
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states [20, 33]. Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
We consider QED in the presence of a coherent background EM field and reformulate it in terms of coherent-field QED. To this end, we assume that a coherent background laser field is always present. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
@@
corresponding to a background field description prior to imposing fixed boundary conditions.
@@
Under this preparation, the Schrödinger equation governing the “coherent laser field plus other particles” state is
$i\hbar\partial_{t}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ $=\hat{H}_{\mathrm{S}}|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle,$ (12)
@@
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) [15, 2] evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture [20, 33], the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
@@
with the formal solution
@@
Thus the time evolution of the quantum state $|\Psi_{\mathrm{S}}(t)\rangle$, with the coherent laser field factored out, incorporates the influence of the laser EM field ($\mathcal{A}$) through its explicit appearance in the evolution operator.
This confirms that the Schrödinger picture evolution operator consistently incorporates the coherent background field through displacement, while maintaining full quantum dynamics for all other degrees of freedom. In this way, the framework provides a rigorous starting point for connecting the Schrödinger picture with the Heisenberg picture in the presence of a coherent laser background.
The key point of this subsection is that the usual formulation of QED in prescribed background fields emerges directly from the coherent-field QED framework once the coherent state is factored out in the Schrödinger picture. In particular, starting from the fully quantum description $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$, the displacement by a coherent state produces the familiar fixed-background description $|\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] widely used in strong-field QED. This shows that the fixed-background field approximation is not the starting point but rather a derived limit within coherent-field QED. Moreover, because depletion and path-integral formulations require transitions between different coherent states, the fundamental structure of this work is the underlying coherent-field QED structure defined by
$\big(\,|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle,\,\,$ $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]\,\big),$ (22)
not its fixed-background reduction.